### v1

# Zero-Shot Referring Expression Comprehension via Visual-Language True/False Verification

Abstract
Referring Expression Comprehension (REC) is usually addressed with task-trained grounding models. We show that a zero-shot workflow, without any REC-specific training, can achieve competitive or superior performance. Our approach reformulates REC as box-wise visual–language verification: given proposals from a COCO-clean generic detector (YOLO-World), a general-purpose VLM independently answers True/False queries for each region. This simple procedure reduces cross-box interference, supports abstention and multiple matches, and requires no fine-tuning. On RefCOCO, RefCOCO+, and RefCOCOg, our method not only surpasses a zero-shot GroundingDINO baseline but also exceeds reported results for GroundingDINO trained on REC and GroundingDINO+CRG. Controlled studies with identical proposals confirm that verification significantly outperforms selection-based prompting, and results hold with open VLMs. Overall, we show that workflow design, rather than task-specific pretraining, drives strong zero-shot REC performance.
Index Terms:
Referring expression comprehension, large language modeling inference, zero-shot reasoning
I Introduction
Figure 1: The concept behind our verification-first workflow, which enables outstanding accuracy in a training-free zero-shot setting.
Referring Expression Comprehension (REC) [1] is the task of locating the specific object described by a natural language expression (e.g., ‘the small red mug on the left’) in a given image. REC supports interactive perception and grounding for applications such as camera auto-directing and framing in live streaming, video content editing, and voice-directed robotic/device control. The standard output is a single bounding box that identifies the referent.
Despite strong progress in large vision–language models (VLMs), e.g., CLIP[2], LLaVA[3], GPT-4o[4], REC remains difficult. The task demands fine-grained language understanding and grounded visual reasoning to select one instance among many look-alike objects. Descriptions blend attributes, parts, and spatial/ordinal cues, often in scenes with occlusion and scale changes, so that close similarities are common. Furthermore, off-the-shelf VLMs without REC training rarely produce accurate bounding boxes as they are not calibrated for instance-level localization.
In the supervised setup, models learn the REC mapping directly from annotated referring datasets. Approaches include end-to-end grounding architectures and cross-modal matching heads that predict a box from an image–text pair. Text-conditioned detectors such as GroundingDINO[5] show substantial gains when trained with RefCOCO-style supervision. Supervised large multimodal models, including CogVLM[13] variants finetuned for grounding, currently define the-state-of-the-art under this regime.
Another common setup uses grounding-trained proposal generators with inference-time decision logic. A text-conditioned detector (often GroundingDINO) produces language-aware region proposals with bounding boxes, and a separate module refines or selects among them at test time. CRG[6] is a representative example that re-scores proposals with a VLM. Although the final stage operates without additional training, the overall pipeline is not zero-shot at the system level, because the proposal source has been trained on REC-style data.
Finally, in zero-shot workflows[10][11], both components are off-the-shelf: a generic detector supplies candidate boxes, and a VLM provides the decision logic via prompting. Neither of the models is trained on REC. The decision logic is commonly based on global ranking of the top-K candidates. Performance is bounded by proposal recall of the generic detector. The workflow, including prompt design, decomposition into attributes and relations, and tie-breaking, becomes the primary lever for accuracy and robustness. Our method targets this zero-shot workflow regime.
Our method treats REC as box-wise verification on the original image. A generic detector proposes candidates. For each proposal, we render the image with only that box drawn and ask a VLM a single binary question: does this box match the description? The VLM returns True/False without scores. This first pass shrinks the candidate set. If exactly one box is True, we return it. If multiple boxes are True, we render the image with only those True boxes overlaid together and ask the VLM to select the one that best fits the description, with all False boxes being excluded from this step. If all boxes are False, we fall back to a global selection prompt over the full proposal set.
On RefCOCO[8], RefCOCO+, and RefCOCOg[9], our verification-first workflow achieves ACC@0.5 scores of 81.2%, 76.3%, and 75.9%, respectively. This marks a substantial improvement over the zero-shot GroundingDINO baseline without RefCOCO training (50.4%, 51.4%, 60.4%). Our approach sets a new state of the art for these benchmarks under zero-shot conditions. Here, ‘zero-shot’ means that no model is trained on the REC task, the detector is COCO-clean (trained without COCO[7] data), and the VLM is used strictly off-the-shelf—though its pretraining corpus may include COCO images, which is not explicitly known. It also surpasses the reported performance of GroundingDINO with RefCOCO training by 11.3% on average. Compared with a single-shot selection approach that uses hand-crafted prompts over the same proposals and VLMs, our proposed True/False verification workflow improves ACC@0.5(%) by 18.0% on average across the three benchmarks and two VLMs.
Zero-shot REC is valuable for both practicality and scientific insight. Practically, it allows developers to build solutions from widely available, off-the-shelf components instead of relying on task-specific models that require curated labels, specialized training, and repeated re-tuning as domains shift. More importantly, zero-shot REC offers a clean testbed for workflow design: can a carefully constructed procedure with generic models to solve a specialized task? Conceptually, REC decomposes into two basic capabilities: (1) visual–language comprehension and reasoning, and (2) deriving a bounding box for a given region. If these fundamentals can be combined to solve REC, it suggests that other composite tasks may likewise be addressed modularly by reusing general-purpose models for their constituent subtasks.
Our results echo the trends in LLMs where prompt design, tool use, and staged reasoning often replace fine-tuning. By showing that True/False verification narrows and often resolves the decision problem, we present a template that extends beyond REC to other complex vision tasks (or even generic LLM based tasks). Our central insight includes:
  • Verification over comparison. Modern vision–language models reason more reliably when prompted with a single, concrete hypothesis rather than a comparative choice among many alternatives. Each query of box-wise verification isolates a single candidate and removes cross-candidate coupling, order effects, and prompt entanglement, thereby minimizing interference.
  • Selection as atomic checks. Any multi-candidate selection can be expressed as a set of independent True/False tests. Running verification once per proposal converts a joint decision into parallel atomic decisions and yields a pruned candidate set. Even in the multi-True case, this reduction sharply shrinks the search space and empirically improves the final selection’s accuracy and stability.
II Method
TABLE I: Zero-shot, verification-first REC
1. Input image $I$, description $s$, detector $D$, VLM $F$.
2. $c\leftarrow F$ infer class from $s$.
3. $\mathcal{B}\leftarrow D(I,c)$ (candidate boxes).
4. For each $b_{i}\in\mathcal{B}$: overlay $\tilde{I}_{i}$ and query $F(\tilde{I}_{i},s)\!\to\!y_{i}\in\{\texttt{True},\texttt{False}\}$.
5. $\mathcal{T}\leftarrow\{i\mid y_{i}=\texttt{True}\}$. If $|\mathcal{T}|=1$, return that box.
6. If $|\mathcal{T}|>1$, overlay & index only $\{b_{i}\}_{i\in\mathcal{T}}$; ask $F$ to pick best; return it.
7. If $\mathcal{T}=\varnothing$, overlay & index all boxes; ask $F$ to pick best; if “none,” abstain.
II-A Problem Setup and Zero-Shot Setting
REC takes an image and a description and returns the best-matching bounding box (or abstains if none fits). We use a zero-shot workflow with only off-the-shelf models: a basic, class-conditioned detector without grounding capability to propose boxes, and a general-purpose VLM to verify them. [^1] Zero-shot means no REC training; the detector is COCO-clean; the VLM is used as-is and may or may not have seen COCO during pretraining. This shifts task specialization from model weights to the procedure.
II-B Verification-First Workflow
Our workflow reformulates REC as box-wise visual–language verification and resolves the final choice with a lightweight VLM-driven selector. As shown in Figure 1 and Table I, the workflow has the following steps:
  1. (a)
    Class identification. From the natural-language description, ask the VLM to name the most relevant object class (e.g., person, dog, car). Use this class to focus the detector.
  2. (b)
    Class-conditioned proposals. Run the detector on the image for that class to produce a set of candidate bounding boxes.
  3. (c)
    Box-wise verification. For each candidate box, draw it on the image and ask the VLM: “Does the description apply to this box?” Record a True/False judgment for each box.
  4. (d)
    Decision rule (with abstention).
    • If exactly one box is labeled True, output that box.
    • If multiple boxes are labeled True, show only those boxes (with indices) and ask the VLM to choose the best-fitting one; output it.
    • If all boxes are labeled False, show all boxes (with indices) and ask the VLM to choose the best match; if it replies “none,” abstain.
(a) $p$ threshold vs. $q_{1}$ and $q_{2}$.
(b) $p$ threshold as a function of $q$.
(c) Accuracy gain by verification vs selection.
Figure 2: Threshold and gain analyses derived from the two-candidate model (Sec. III).
TABLE II
REC benchmarking (ACC@0.5 (%)).
Regime Method RefCOCO RefCOCO+ RefCOCOg
Val TestA TestB Val TestA TestB Val Test
Supervised CogVLM (REC-trained) 92.6 94.3 91.5 85.2 89.6 79.8 88.7 89.4
Supervised GroundingDINO (REC-trained) 77.3 72.5 72.0 59.3 66.3
Workflow with REC GroundingDINO + CRG (REC-trained rerank) 81.6 73.2 77.0 60.0 69.6
Strict zero-shot GroundingDINO (zero-shot baseline) 50.4 57.2 43.2 51.4 57.6 45.8 60.4 59.5
Zero-shot VLM (baseline) GPT-4o (organic / vanilla selection) 8.4 7.3 9.1
Zero-shot on REC (ours) Selection (LLaVA, single-shot) 34.7 34.6 44.4
Zero-shot on REC (ours) Verification-first (LLaVA) 44.6 42.3 50.7
Zero-shot on REC (ours) Selection (GPT-4o, single-shot) 70.1 73.5 65.4 66.7 68.3 60.2 69.7 68.4
Zero-shot on REC (ours) Selection (GPT-4o, majority voting) 71.7 67.4 69.9
Zero-shot on REC (ours) Verification-first (GPT-4o) 81.2 86.2 74.5 76.3 81.2 69.4 75.9 74.6
Rather than asking the VLM to choose one box among many, we verify each candidate independently with a binary True/False query. Although the binary verification workflow may appear simple and superficially similar to selection-based prompting, it behaves quite differently in practice. We expect the performance improvements to stem from three mechanisms:
  • Reduced cross-box interference: Binary verification workflow decouples candidates and reduces cross-box interference, especially when the detector generates many proposals. Focusing the model on a single highlighted region concentrates its reasoning and limits distraction from nearby objects.
  • Built-in error control: Binary verification also provides a simple yet strong error-control signal: when exactly one box is labeled True, without being told how many positives to expect, the VLM has singled out a unique region that satisfies the description. This outcome is far more likely to reflect genuine comprehension and reasoning than chance.
  • Pruned candidates: When multiple boxes are True, the set of candidates is already pruned, allowing a lightweight second-stage tie-break over a smaller and cleaner pool.
III Analysis
Consider two proposals: bounding box 1 is correct and bounding box 2 is a distractor. Under a selection prompt, let us define $p$ as the probability of selecting bounding box 1, so
$A_{\text{sel}}=p.$ (1)
Under box-wise True/False verification, let us define $q_{1}$ and $q_{2}$ as probabilities of bounding box 1 and bounding box 2 being labeled True, respectively. When both are True or both are False, it falls back to the selection setup. Then the accuracy of the verification approach is
$A_{\text{ver}}=q_{1}(1-q_{2})+q_{1}q_{2}p+(1-q_{1})(1-q_{2})p.$ (2)
To determine when selection outperforms verification ($A_{\text{sel}}\geq A_{\text{ver}}$), we can have:
$p\;\geq\;1\;-\;\frac{q_{2}(1-q_{1})}{\,q_{1}(1-q_{2})+q_{2}(1-q_{1})\,}.$ (3)
Thus selection must exceed the threshold in (3) to beat verification.
As shown in Figire 2 (a), for any fixed $q_{2}<0.5$, the threshold of $p$ is an increasing and concave function of $q_{1}$ and lies above the identity line $p=q_{1}$ for $0<q_{1}<1$. Hence, to match the accuracy of verification, the selection method must achieve $p>q_{1}$ (i.e., be more reliable than the verifier’s true-positive rate). Design-wise, we aim to keep $q_{2}<0.5$. Note that $q_{2}=0.5$ corresponds to random guessing on the distractor, so effective prompting/overlays should push $q_{2}$ well below this. With a reasonably capable VLM and small $q_{2}$, the verification succeeds more easily, and selection must be correspondingly stronger to keep up.
To further simplify, assume $q_{1}=q$ and $q_{2}=1-q$. As shown in Figure 2 (b), $p$ threshold lies above the identity line $p=q$ for $q>0.5$. To match verification, the selection method must achieve a probability $p$ strictly higher than the verifier’s true-positive rate $q$. Figure 2 (c) plots the gap between $p$ threshold and $q$, which peaks at about $0.145$ near $q\!\approx\!0.7$. In other words, when the verifier labels the correct box True with probability $q\!\approx\!0.7$ (and the distractor with $1-q\!\approx\!0.3$), selection would need $p\!\approx\!0.845$—roughly a +0.15+0.15 absolute increase—to achieve the same accuracy.
One might view this comparison as unfair: the selection baseline uses a single run, whereas our verification-first pipeline uses at least two rounds (and a third only to break ties). In principle, we could also run the selection procedure three times and take a majority vote; under an i.i.d. assumption this yields an effective accuracy of $3q^{2}-2q^{3}$ and would require us to compare against the condition $p>3q^{2}-2q^{3}$, which is stricter than the threshold in Eq. (3). However, two caveats matter. First, majority-vote selection requires three rounds, while for large $q$ the verification-first scheme typically concludes in two; thus the compute budgets are not apples-to-apples. Second, repeated calls to the same VLM on the same instance are rarely i.i.d. — especially at the low temperatures commonly used for selection — so multiple runs often reproduce the same output, making majority voting little better than a single run. (This assumption is validated in Table 2). By contrast, verification-first queries different regions/objects across rounds, making the i.i.d. assumption more plausible. For these reasons, we evaluate the selection baseline with a single run in this analysis.
The two-box analysis omits two sources of advantage in practice. First, cross-box interference is suppressed under verification because each query presents a single highlighted region. This typically raises the verifier’s true-positive rate while not improving the selector, widening the gap required for selection to match verification. Second, with more than two proposals, verification induces pruning, while the two-box model cannot reflect the gain as there is no meaningful shrinkage when $n=2$.
IV Experiments
We evaluate on RefCOCO, RefCOCO+, and RefCOCOg, i.e., the standard REC benchmarks derived from MS-COCO, each providing image-expression pairs. Following RefCOCO conventions, a detection is correct if its Intersection-over-Union (IoU) with the ground-truth box exceeds 0.5. We therefore report ACC@0.5. We consider the following approaches in our experiments for comparison:
  • CogVLM (supervised). A vision–language model trained with REC supervision and evaluated in the standard supervised regime.
  • GroundingDINO (REC-trained). GroundingDINO fine-tuned on REC datasets to directly localize the referred object.
  • GroundingDINO + CRG (REC-trained workflow). A re-ranking workflow that scores proposals from REC-trained GroundingDINO using language/context cues to select the final box.
  • GroundingDINO (zero-shot baseline). A pretrained text-conditioned detector applied without any REC fine-tuning to output the top-1 box.
  • GPT-4o (vanilla REC). Off-the-shelf GPT-4o is shown the image with description and asked to generate the best-matching bounding box in a single prompt (no REC training).
  • Verification-first (LLaVA, ours). Use YOLO-world[14], a COCO-clean basic class-conditioned detector (no grounding), proposes boxes, and off-the-shelf and open LLaVA verifies each box with a True/False query, with abstention/tie-breaks as needed (no REC training).
  • Verification-first (GPT-4o, ours). Same verification pipeline as above but using GPT-4o instead of LLaVA.
  • Selection (LLaVA, ours). Control variant that uses the same proposals and an open VLM (LLaVA) issues a single-shot selection prompt rather than box-wise verification.
  • Selection (GPT-4o, single-shot, ours). Same selection pipeline as above but using GPT-4o instead of LLaVA.
  • Selection (GPT-4o, majority voting, ours). Same selection pipeline as above, using GPT-4o, but using majority voting of three runs. Temperature is configured to 1.0.
We use YOLO-World as a basic class-conditioned detector (no grounding), trained without COCO for dataset cleanliness. This COCO-clean choice comes with a modest cost, i.e., approximately 10 percentage points mAP@0.5 lower than comparable COCO-trained YOLO variants on detection tasks. For VLMs, we evaluate GPT-4o and LLaVA-vicuna-13b off-the-shelf. A “organic” GPT-4o baseline attains less than 10% ACC@0.5 on these benchmarks, suggesting that it is not REC-trained. The strong numbers we obtain stem from the workflow rather than hidden task-specific pretraining.
As shown in Table 2, except for the fully supervised CogVLM, our verification-first (GPT-4o) approach outperforms all other approaches across RefCOCO/RefCOCO+/RefCOCOg. Using the same detector, VLM, and proposal sets, verification consistently beats selection, including the selection with majority voting, by about 6 to 10 absolute points across most of the splits, underscoring that workflow—not just model choice—drives the gains. We also observe that GPT-4o surpasses LLaVA in both verification and selection, indicating that VLM capability is critical in zero-shot REC. In summary: (i) a carefully designed workflow, like the True/False verification workflow, can deliver large improvements even with identical components, and (ii) fundamental models (a basic, non-grounding detector and a general-purpose VLM) can be composed to solve a composite task like REC without task-specific training—though the strength of each component still matters.
V Conclusion
We present a verification-first, zero-shot REC workflow that achieves high performance. Extensive experiments show consistent gains over single-shot selection and over REC-trained GroundingDINO(+CRG). The results indicate workflow design, rather than task-specific pretraining, drives the improvement. Future work will extend to other vision tasks.